# Understanding Co-speech Gestures in-the-wild

### v1

@@ 6.2 Impact of loss functions
Loss Retrieval Spotting ASD R@5 $\uparrow$ R@10 $\uparrow$ MR $\downarrow$ Acc. $\uparrow$ Acc. $\uparrow$ Seq. contrastive 11.60 21.40 43 18.83 42.8 Word coupling 7.90 13.40 78 51.98 19.6 Seq. + Word coupling 17.40 28.40 22 63.00 59.0
@@ 6.3 Fusion techniques
In Table 7, we ablate different ways to fuse the speech and text features. The first case is to not fuse at all and have two separate pairwise contrastive losses: gesture-audio and gesture-text. We can see in the first two rows that this is an inferior design. In fact, it is better to train with a single contrastive head after fusing the speech and text embeddings (rows $3$, $4$). The fusion strategy of choice would be to concatenate, rather than average.
@@
Retrieval Spotting ASD R@5 $\uparrow$ R@10 $\uparrow$ MR $\downarrow$ Acc. $\uparrow$ Acc. $\uparrow$ Pairwise (with text) 9.39 15.58 70 34.31 29.6 Pairwise (with audio) 9.80 16.60 72 23.67 31.4 Late fusion (avg.) 16.00 24.60 30 57.27 54.0 Late fusion (concat.) 17.40 28.40 22 63.00 59.0
@@ 7 Conclusion
In this work, we learn a joint embedding space that captures cross-modal relationships with gestures, speech, and language. We show that we can learn such an embedding space with weak supervision using a careful design of two loss functions. We evaluate these new representations on three new downstream tasks and manually curated test sets. We observe that the two modalities, i.e., speech and text, learn complementary features that can be useful for different kinds of gesture-related tasks. One promising future direction would be to explore 2D and 3D keypoint based inputs to make the network computationally lighter and less susceptible to distracting features.
@@ Appendix A Most Gestured Words
Appendix B Model Details
In Table 8, we provide detailed description of the model architecture. We will release the code and models to support future research.
@@ Appendix B Model Details
Appendix C Dataset Visualization
In Figure 6, we present examples from our manually annotated AVS-Spot test set (curated from the publicly available AVSpeech test dataset [20]), designed to evaluate downstream gesture spotting performance. As shown, the dataset includes a diverse collection of unique words, carefully curated to ensure clear and contextually appropriate gestures. For instance, in row-$1$, the word “little” is accompanied by a gesture where two fingers move close together to indicate a small size; in row-$2$, the speaker points backward to represent the word “back”; and in row-$6$, an inward motion is used to indicate “my”.
@@ Appendix C Dataset Visualization
Appendix D Qualitative Results
In Figure 7, we show additional qualitative examples for gesture spotting. In the left text panel, the red-highlighted word represents the keyword to be spotted, as curated in the AVS-Spot test set. The word-labeled vertical columns, separated by yellow lines, indicate the word boundaries derived from speech-text alignment. JEGAL successfully spots most of these keywords, as shown by the red heatmaps. Notably, the boundaries may vary slightly since speakers often gesture and speak at slightly different times, highlighting the inherent challenges of our weakly-supervised gesture representation learning task.
In Figure 8, we present additional examples demonstrating that audio-based gesture spotting tends to focus on “stressed regions” in speech, unlike text-based spotting. This difference is evident from the audio and text heatmaps for each sample. In Figure 8, our model detects the stressed keywords “every” and “respond”, whereas the text-only model misses these words. Evidently, the audio-only model looks for word emphasis cues (indicated by high pitch) as such words are more likely to be gestured. This would be difficult to infer from text modality alone. These examples illustrate the advantages of leveraging audio cues for gesture spotting.
Appendix E Limitations and Areas of Improvement
Our work is the first to tackle large-scale co-speech gesture understanding. We highlight some of the limitations of our approach here. Having trained predominantly on TED/TEDx talks, our models can struggle with co-speech conversational gestures. Further, the TED/TEDx talk setting inherently encourages continuous gesturing and moving the arms, which might not always be the case for other settings (like conversational interactions). Another aspect the model struggles with is when there are limited gesture actions or hand movements that are unrelated to speech. Finally, given that we learn with only weak sequence-level supervision, the model can “find shortcuts” by focusing on simple rhythmic hand movements that occur in certain gestures classes like the beat gestures. This can affect the representation quality of iconic and deitic gestures that contain clear semantic meaning. While we still show that our models can spot such gestures, future works can focus on improving this imbalance in gesture classes.
Appendix F Potential Negative Societal Impacts
While our research significantly contributes to advancing gesture understanding, there are some potential risks of surveillance, as the system could infer conversations from a distance by identifying words/phrases. Nonetheless, we believe the benefits outweigh these risks, as the technology enhances human-machine interaction by integrating non-verbal cues. According to the $55$% rule [1], non-verbal communication constitutes $55$% of overall communication. This highlights the importance of enabling machines to engage in holistic, natural interactions with humans by understanding non-verbal elements like gestures.
Figure 7: Additional gestured word spotting results on AVS-Spot dataset. Keywords are highlighted in red on the left panel and the speech-based force alignment word boundaries are marked by yellow lines. JEGAL successfully spots the gestured keywords, demonstrating its robustness across diverse gestures and speakers. The red triangles zoom into the corresponding frames where JEGAL detects the keywords, clearly aligning with the gestures. Note that in some cases (e.g., rows $2$ and $3$), ground-truth boundaries may slightly differ, as the speaker can gesture and utter the same word at slightly different times. JEGAL effectively estimates the approximate intervals where the target word is gestured.
Figure 8: Examples highlighting the role of stressed speech regions in audio-based gesture spotting. The audio-only model successfully detects the stressed keywords “every” and “respond”, whereas the text-only model misses these words. Evidently, the audio-only model looks for word emphasis cues (indicated by high pitch) as such words are more likely to be gestured. This would be difficult to infer from text modality alone. These examples illustrate the advantages of leveraging audio cues for gesture spotting.

### v2

@@ 6.2 Impact of loss functions
Loss Retrieval Spotting ASD R@5 $\uparrow$ R@10 $\uparrow$ MR $\downarrow$ Acc. $\uparrow$ Acc. $\uparrow$ Seq. contrastive 12.20 23.60 45 20.83 44.2 Word coupling 8.50 14.60 76 52.46 14.8 Seq. + Word coupling 18.80 30.80 31 63.60 48.0
@@ 6.3 Fusion techniques
In Table 7, we ablate different ways to fuse the speech and text features. The first case is to not fuse at all and have two separate pairwise contrastive losses: gesture-audio and gesture-text. We can see in the first two rows that this is an inferior design. In fact, it is better to train with a single contrastive head after fusing the speech and text embeddings (rows $3$, 44). The fusion strategy of choice would be to concatenate, rather than average.
@@
Retrieval Spotting ASD R@5 $\uparrow$ R@10 $\uparrow$ MR $\downarrow$ Acc. $\uparrow$ Acc. $\uparrow$ Pairwise (with text) 9.39 15.58 70 34.31 29.6 Pairwise (with audio) 9.80 16.60 72 23.67 31.4 Late fusion (avg.) 17.00 26.40 40 56.04 41.2 Late fusion (concat.) 18.80 30.80 31 63.60 48.0
@@ 7 Conclusion
In this work, we learn a joint embedding space that captures cross-modal relationships with gestures, speech, and language. We show that we can learn such an embedding space with weak supervision using a careful design of two loss functions. We evaluate these new representations on three new downstream tasks and manually curated test sets. We observe that the two modalities, i.e., speech and text, learn complementary features that can be useful for different kinds of gesture-related tasks. One promising future direction would be to explore 2D and 3D keypoint-based inputs to make the network computationally lighter and less susceptible to distracting features.
@@ Appendix A Most Gestured Words
Appendix B Additional Evaluations and Analysis
B.1 Gesture Word Spotting: Evaluation in challenging conditions
Our evaluation set (constructed from AVSpeech) includes a diverse range of samples: (i) non-frontal videos, (ii) varying lighting conditions, (iii) a wide variety of speakers, and (iv) conversational videos (from which we extract segments featuring a single speaker). In this section, we specifically benchmark the performance of JEGAL on these challenging subsets. We label the AVS-Spot dataset with new metadata: (i) lighting conditions (dim, medium, bright), and (ii) speaker poses (frontal vs. non-frontal). Fig 6 illustrates the diversity of the test set.
Table 8 reports the spotting accuracy across these subsets. We find that the model performs best on brightly lit videos, with similar accuracy for dim and medium lighting.
Table 8: Evaluation in challenging conditions: JEGAL outperforms prior models in all settings.
Method Lighting Speaker pose
Dim Medium Bright Frontal Non-frontal
GestSync [23] 9.67 22.92 8.33 21.59 17.80
GestDiffuClip [8] 15.6 19.7 21.8 19.35 19.90
JEGAL (Ours) 61.29 62.58 77.77 62.76 68.49
Figure 6: The AVS-Spot test set is quite diverse – some examples are shown above. Additionally, we annotate the clips in AVS-Spot for frontal/non-frontal views and lighting and analyze the performance on these individual subsets.
B.2 Effect of Modality dropping
We present the impact of dropping text and audio modalities at varying rates on the spotting task in Table 9. A drop rate of $30\%$ means that during training, either text or audio is randomly dropped in $30\%$ of the batch samples. Dropping the modalities at $50\%$ performs the best across all inference-time settings.
Table 9: Dropping modalities evenly during training works best.
   Drop %    Accuracy $\uparrow$
   T    A    TA
   30%    52.2    38.6    63.2
   50% (JEGAL)    61.0    41.8    63.6
   70%    61.3    42.2    62.6
B.3 Computational efficiency
Table 10 shows the inference time (averaged across ten runs) for a $5$-second input on a single NVIDIA $V100$ GPU. Our model can process $\approx 52$ frames per second, indicating that the inference is quite fast but is not streaming-capable yet, as the bidirectional transformer attends to all future frames provided as context.
Table 10: Inference time analysis for 5-second input.
Visual Enc. Text Enc. Audio Enc. Total
inf. time (sec) $\downarrow$ 1.84 0.18 0.07 2.09
B.4 Where does the model focus on?
We visualize the activation maps of the visual features of JEGAL to see which spatial region of the video the model focuses on. In Fig 7, we see that the model focuses on the hand gestures.
Figure 7: We plot the activation maps of the visual features of JEGAL. We can see that JEGAL focuses strongly on the hand gestures.
Appendix C Model Details
In Table 11, we provide detailed description of the model architecture. The code and models have been released to support future research.
@@ Appendix C Model Details
Appendix D Dataset Visualization
In Figure 8, we present examples from our manually annotated AVS-Spot test set (curated from the publicly available AVSpeech test dataset [18]), designed to evaluate downstream gesture spotting performance. As shown, the dataset includes a diverse collection of unique words, carefully curated to ensure clear and contextually appropriate gestures. For instance, in row-$1$, the word “little” is accompanied by a gesture where two fingers move close together to indicate a small size; in row-$4$, the speaker points backward to represent the word “back”; and in row-$6$, the fingers of both hands move in a distinctive pattern to indicate “hashtag”.
Appendix E Qualitative Results
In Fig 9, we show additional qualitative examples for gesture spotting. In the left text panel, the red-highlighted word represents the keyword to be spotted, as curated in the AVS-Spot test set. The word-labeled vertical columns, separated by yellow lines, indicate the word boundaries derived from speech-text alignment. JEGAL successfully spots most of these keywords, as shown by the red heatmaps. Notably, the boundaries may vary slightly since speakers often gesture and speak at slightly different times, highlighting the inherent challenges of our weakly-supervised gesture representation learning task.
In Fig 10, we present additional examples demonstrating that audio-based gesture spotting tends to focus on “stressed regions” in speech, unlike text-based spotting. This difference is evident from the audio and text heatmaps for each sample. In Fig 10, our model detects the stressed keywords “specific” and “respond”, whereas the text-only model misses these words. Evidently, the audio-only model looks for word emphasis cues (indicated by high pitch) as such words are more likely to be gestured. This would be difficult to infer from text modality alone. These examples illustrate the advantages of leveraging audio cues for gesture spotting.
Appendix F Limitations and Areas of Improvement
Our work is the first to tackle large-scale co-speech gesture understanding. We highlight some of the limitations of our approach here. One aspect the model struggles with is when there are limited gesture actions or hand movements that are unrelated to speech. Another shortcoming is that since we learn with only weak sequence-level supervision, the model can “find shortcuts” by focusing on simple rhythmic hand movements that occur in certain gestures classes like the beat gestures. This can affect the representation quality of iconic and deitic gestures that contain clear semantic meaning. While we still show that our models can spot such gestures, future works can focus on improving this imbalance in gesture classes.
Appendix G Potential Negative Societal Impacts
While our research significantly contributes to advancing gesture understanding, there are some potential risks of surveillance, as the system could infer conversations from a distance by identifying words/phrases. Nonetheless, we believe the benefits outweigh these risks, as the technology enhances human-machine interaction by integrating non-verbal cues. According to the 5555% rule [1], non-verbal communication constitutes 5555% of overall communication. This highlights the importance of enabling machines to engage in holistic, natural interactions with humans by understanding non-verbal elements like gestures.
@@ Appendix G Potential Negative Societal Impacts
Figure 9: Additional gestured word spotting results on AVS-Spot dataset. Keywords are highlighted in red on the left panel and the speech-based force alignment word boundaries are marked by yellow lines. JEGAL successfully spots the gestured keywords, demonstrating its robustness across diverse gestures and speakers. The red triangles zoom into the corresponding frames where JEGAL detects the keywords, clearly aligning with the gestures. Note that in some cases (e.g., rows 22 and 44), ground-truth boundaries may slightly differ, as the speaker can gesture and utter the same word at slightly different times. JEGAL effectively estimates the approximate intervals where the target word is gestured.
Figure 10: Examples highlighting the role of stressed speech regions in audio-based gesture spotting. The audio-only model successfully detects the stressed keywords “specific” and “respond”, whereas the text-only model misses these words. Evidently, the audio-only model looks for word emphasis cues (indicated by high pitch) as such words are more likely to be gestured. This would be difficult to infer from text modality alone. These examples illustrate the advantages of leveraging audio cues for gesture spotting.